How far can a man drive out into the country at the average rate of 60 miles per hour and return over the same road at the average rate of 45 miles per hour if he travels a total of 7 hours?
Answer by checkley77(12844)

60X=45(7-X)
60X=315-45X
60X+45X=315
105X=315
X=3 HOURS IS THE TIME FOR THE 60MPH DRIVE
7-4=3 HOURS IS THE TIME FOR THE RETURN TRIP.
PROOF:
60*3=45*4
180=180